### v1

# GOOD: Towards Domain Generalized Orientated Object Detection

Abstract
Oriented object detection has been rapidly developed in the past few years, but most of these methods assume the training and testing images are under the same statistical distribution, which is far from reality. In this paper, we propose the task of domain generalized oriented object detection, which intends to explore the generalization of oriented object detectors on arbitrary unseen target domains. Learning domain generalized oriented object detectors is particularly challenging, as the cross-domain style variation not only negatively impacts the content representation, but also leads to unreliable orientation predictions. To address these challenges, we propose a generalized oriented object detector (GOOD). After style hallucination by the emerging contrastive language-image pre-training (CLIP), it consists of two key components, namely, rotation-aware content consistency learning (RAC) and style consistency learning (SEC). The proposed RAC allows the oriented object detector to learn stable orientation representation from style-diversified samples. The proposed SEC further stabilizes the generalization ability of content representation from different image styles. Extensive experiments on multiple cross-domain settings show the state-of-the-art performance of GOOD. Source code will be publicly available.
1 Introduction
Oriented object detection occupies a unique intersection between computer vision and remote sensing [47, 12, 51, 53], involving the prediction of oriented bounding boxes for aerial objects. Given the bird’s-eye view perspective of an image sensor [56, 55, 3, 4, 2, 62], aerial objects pose arbitrary orientations. Therefore, in addition to the conventional task of accurately identifying and localizing objects [40, 30, 27], an oriented object detector must also provide precise predictions of an object’s rotation angle [35, 49, 50, 54].
Existing oriented object detectors typically rely on the assumption that the source domain during training and the target domain during inference share the same statistical distribution. However, this assumption often diverges from reality. As shown in Fig. 1a, aerial images from different domains have great style variation, caused by illumination, imaging sensor, landscape and etc. A pre-trained oriented detector on Google Earth images from American landscape (e.g., SODA [8] as source domain) need to conduct inference on numerous unseen target domains, e.g., Gaofen-2 images from Chinese landscape in DOTA [47] domain (Fig. 1b). Developing a domain generalized oriented object detector enables the inference across diverse, previously unobserved target domains. Regrettably, as far as our knowledge extends, this area remains largely unexplored.
Figure 1: (a) Oriented object images different domains (e.g., SODA [8], DOTA [47], FAIR1M [39], HRSC [31]) can have dramatic style variation (e.g., spatial resolution, contrast, illumination), which leads to different feature distribution; (b) Domain generalized oriented object detection aims to learn an oriented object detector from only source domain that can be well generalized to arbitrary unseen target domains.
In this paper, we make an early exploration to learn domain generalized oriented object detectors. The great style variation between the source domain and unseen target domains is usually reflected from the image appearances that have different color distribution (e.g., intensity, contrast) [22, 9, 33, 5], which are the input of an oriented object detector pre-trained on a certain source domain. The different color distribution makes the pre-trained oriented object detector difficult to perceive the activation patterns and to build the feature representation as learnt on the source domain [21, 9]. The shift of activation patterns on the target domains inevitably poses two key challenges for oriented object detectors.
First of all, the shift of activation patterns accumulates from shallow to deep of an oriented object detector, and inevitably confuses and degrades the high-level representation, e.g., region of interests (RoIs). Consequently, as illustrated in Fig. 2a, the category-level representation, also known as a kind of content information, can be less reliable for a detector to precisely identify an object. A same object category from different styles can be difficult to cluster into the same category. Secondly, such statistic variation from the target domain can also lead to less representative orientation information (in Fig. 2b). This is because the orientation is usually regressed from the RoIs. As the RoIs degrade on the target domains, the orientation inevitably becomes less precise. As illustrated in Fig. 2c, the orientation of the prediction can be quite different from the ground truth.
Figure 2: Two key challenges when an oriented object detector generalizes to unseen target domains. (a) degraded category prediction: multiple ships are not detected; (b) imprecise oriented prediction: the rotation angle does not align with the object orientation. (c) Analysis from feature space for both challenges. Oriented object detector: ReDet [16]; Source domain: DOTA [47]; Unseen target domain: FAIR1M [39] & HRSC [31].
We propose a domain generalized oriented object detector, dubbed as GOOD, to address the above challenges. First, we aim to allow the GOOD to perceive as many as styles when training on the source domain. The emerging contrastive language-image pre-training (CLIP) [36] rests in abundant styles from web-wide images, and provides us a feasible solution. Afterwards, the proposed GOOD consists of two key components, namely, rotation-aware content consistency learning (RAC) and style consistency learning (SEC). The proposed RAC constraints both horizontal region of interests (HRoIs) and rotated region of interests (RRoIs) between the original images and style-hallucinated images, so that both the orientation and content representation is more robust to the cross-domain style variation. On the other hand, the proposed SEC poses a category-wise constraint between the original and style-hallucinated images, which allows the style-hallucinated images to perceive the same content. In this way, the classification head yields more discriminative category prediction on the unseen target domains.
Extensive domain generalization settings on existing oriented object detection benchmarks (i.e., FAIR1M [39], DOTA [47], SODA [8], HRSC [31]) have been designed to benchmark domain generalized oriented object detection. Following prior driving-scene object detection works [46, 29, 42], the same object categories among each dataset are used for experiments. Specifically, four cross-domain settings, namely, FAIR1M-10 classes, SODA-7classes, DOTA-4classes and HRSC-ship, are designed. The proposed GOOD shows substantial improvement on the multiple baselines on all the generalization settings.
Our contribution can be summarized as below.
  • We make an early exploration to the task of learning domain generalized oriented object detectors. A generalized oriented object detector, dubbed as GOOD, is proposed.
  • With CLIP-driven style hallucination, we propose a rotation-aware consistency (RAC) and a style consistency (SEC) learning to represent generalized oriented objects.
  • The proposed GOOD significantly outperforms existing state-of-the-art oriented object detectors on a variety of generalization settings, by up to 3.29% mAP on DOTA and 3.08% on SODA.
2 Related Work
Oriented Object Detection can be categorized as two-stage and one-stage. Typical works of two-stage oriented object detectors include RoI Transformer [11], ReDet [16], Oriented R-CNN [48] and etc. One-stage oriented object detectors can be further categorized into three types. The first type (e.g.,R3Det [51], $S^{2}$-Net [15]) focus on feature alignment between the regression branch and classification branch. The second type focus on angle loss regression [35, 49, 50, 53, 54]. The third type follows the RepPoint paradigm, which leverages the point prediction to bound an oriented object [19, 26]. To learn domain generalized oriented object detectors, we choose the two-stage paradigm not only because it provides a more precise description between orientation and region of interests, but also it provides more access to enrich the style and constraint the content consistency from shallow to deep.
Domain Generalized Object Detection is more challenging than domain adaptive object detection [28, 25, 60]. It aims to learn an object detector from only the source domain data that can be generalized to unseen target domains [29]. More recent works focus on learning domain generalized object detectors on scenarios such as driving scenes [46, 29, 42], 3D point cloud [23, 45], UVA image [43] and etc. On the other hand, domain generalization has received increasingly attention. Multiple advanced machine learning methods (e.g., entropy regularization [61], common-specific low-rank decomposition [34], casual matching [32], extrinsic-intrinsic interaction [44], balance invariance [6], batch normalization embeddings [38]) have been proposed to enhance the generalization ability. However, to the best of our knowledge, none of these works focus on learning oriented object detectors.
Vision-language Model Driven Detection Contrastive language-image pre-training (CLIP) [36] has recently become a domaint visual representation learning trend, which utilizes the image-text pairs to learn visual-semantic representation by pre-training [36, 10, 13]. For object detection community, CLIP has significantly improved the visual representation for open vocabulary object detectors [57, 14, 1, 24, 58], which involve detecting objects based on arbitrary textual descriptions. More recently, Single-DGOD [42] utilizes CLIP for domain generalized detectors on driving scenes. However, in this work, CLIP is only used for semantic augmentation, and does not need to handle the varied orientation in the oriented object detection task.
3 Methodology
Figure 3: schema on generating style-hallucinated samples by the proposed CLIP-driven style hallucination.
Figure 4: Framework overview of the proposed GOOD. After CLIP-driven style hallucination (Sec. 3.2), two key components, namely, rotation-aware consistency learning (RAC, in Sec. 3.3) and style consistency (SEC, in Sec. 3.4) learning, are involved. RAC implements content consistency on both horizontal and rotated region of interests (HRoI and RRoI), as presented in Eq. 6 and Eq. 8, respectively. SEC implements style consistency on category-wise representation from both original and style-hallucinated images (in Eq. 10).
3.1 Problem Setup & Framework Overview
Given a source domain $\mathcal{S}$ and multiple unseen target domains $\mathcal{T}_{1}$, $\cdots$, $\mathcal{T}_{k}$, $\cdots$. In general, the source domain and unseen target domains share the same label space but are under different distributions. For a certain domain $\mathcal{S}$, the joint image and label pair is denoted as $\{\mathbf{I}_{n}^{(\mathcal{S})},(x_{n}^{(\mathcal{S})},y_{n}^{(\mathcal{S})},w_{n}^{(\mathcal{S})},h_{n}^{(\mathcal{S})},\theta_{n}^{(\mathcal{S})})\}_{n=1}^{N_{s}}$, where $n=1,2,\cdots,N_{s}$. Here $x_{n}^{(\mathcal{S})},y_{n}^{(\mathcal{S})},w_{n}^{(\mathcal{S})},h_{n}^{(\mathcal{S})},\theta_{n}^{(\mathcal{S})}$ denote the upper left x-coordinate, upper left y-coordinate, width, height and oriented angle of a oriented region of interest (RRoI), which jointly describe its 2D location in image $\mathbf{I}_{n}^{(\mathcal{S})}$. Also, $N_{s}$ denotes the sample number in domain $\mathcal{S}$.
The learning objective is to learn a oriented object detector $\phi:\mathbf{I}\rightarrow(x,y,w,h,\theta)$ only using the samples from source domain $\mathcal{S}$, and the model $\phi$ is supposed to generalize well on a unseen target domain $\mathcal{T}_{k}=\{(\mathbf{I}_{n}^{(k)})\}_{n=1}^{N_{k}}$.
Fig. 4 gives an overview of the proposed generalized oriented object detector, dubbed as GOOD. After the feature extraction from a backbone, it has two key components, namely, Rotation-aware Consistency Learning (RAC) and Style Consistency Learning (SEC), which are used to decouple the negative impact of different styles on rotation prediction and category prediction, respectively. Overall, it is backbone independent, and can be deployed into off-the-shelf two-stage oriented object detectors that has a step of rotation region of interest (RROI) representation and regression & classification prediction.
3.2 CLIP-driven Style Hallucination
When training on the source domain, a domain generalized oriented object detector is supposed to encounter as many styles as possible. CLIP [36] learns image representation from webwide image-text pairs, which have seen extensive styles in the training stage. The abundant style information could benefit an oriented object detector when generalized to unseen target domains that have unseen styles. To this end, we propose a novel CLIP-driven style hallucination scheme to enrich the style diversity when to learn a generalized oriented object detector.
As an object detector usually leverages features from each stage of a backbone to represent the region of interest, it is necessary to augment the style variation on the features from each stage. Specifically, let $f_{\mathcal{I}}$, $f_{\mathcal{T}}$ and $f$ denote the CLIP image encoder, CLIP text encoder and the backbone of an object detector, respectively. From shallower to deeper, assume $\mathbf{F}_{\mathcal{I}}^{i}\in\mathbb{R}^{W_{i}\times H_{i}\times C_{i}}$ and $\mathbf{F}^{i}\in\mathbb{R}^{W_{i}\times H_{i}\times C_{i}}$ denote the feature representation from the $i^{th}$ block ($i=1,\cdots,4$) of the CLIP image encoder and the detector backbone, respectively, given by
$(\mathbf{F}_{\mathcal{I}}^{1},\mathbf{F}_{\mathcal{I}}^{2},\mathbf{F}_{\mathcal{I}}^{3},\mathbf{F}_{\mathcal{I}}^{4})=f_{\mathcal{I}}(\mathbf{I}),$ (1)
$(\mathbf{F}^{1},\mathbf{F}^{2},\mathbf{F}^{3},\mathbf{F}^{4})=f(\mathbf{I}).$ (2)
Take the feature representation $\mathbf{F}_{\mathcal{I}}^{i}$ from CLIP image encoder as an example. Following existing style-transfer methods [20, 7], the channel-wise normalization mean $\boldsymbol{\mu}_{\mathcal{I}}^{i}$ and standard deviation $\boldsymbol{\sigma}_{\mathcal{I}}^{i}$ carries the style of an image $\mathbf{I}$, computed as
$\boldsymbol{\mu}_{\mathcal{I}}^{i}=\frac{1}{H_{i}W_{i}}\mathbf{F}_{\mathcal{I}}^{i},$ (3)
$\boldsymbol{\sigma}_{\mathcal{I}}^{i}=\sqrt{\frac{1}{H_{i}W_{i}}\sum_{h\in H_{i},w\in W_{i}}(\mathbf{F}_{\mathcal{I}}^{i}-\boldsymbol{\mu}_{\mathcal{I}}^{i})^{2}}.$ (4)
Similarly, let $\boldsymbol{\mu}^{i}$ and $\boldsymbol{\sigma}^{i}$ denote the mean and standard deviation of the feature representation $\mathbf{F}^{i}$ from the detector backbone $f$. Then, the style-diversified counterpart $\mathbf{\widetilde{F}}^{i}$ from $\mathbf{F}^{i}$, which carries the style information from CLIP encoder, can be readily used by AdaIN [20], given by
$\mathbf{\widetilde{F}}^{i}=\boldsymbol{\sigma}_{\mathcal{I}}^{i}\cdot\frac{\mathbf{F}^{i}-\boldsymbol{\mu}^{i}}{\boldsymbol{\sigma}^{i}}+\boldsymbol{\mu}_{\mathcal{I}}^{i}.$ (5)
After that, both the feature representation from the backbone $\mathbf{F}^{i}$ and after style hallucination $\mathbf{\widetilde{F}}^{i}$ ($i=1,\cdots,4$) are processed by a feature pyramid network (FPN) to generate the fused feature representation before and after style hallucination, denoted as $\mathbf{F}$ and $\mathbf{\widetilde{F}}$, respectively.
3.3 Rotation-aware Consistency Learning
Even if plenty of hallucinated styles can be assumed, learning a domain generalized oriented object detector is still particularly challenging, as the orientation information of the content under varied styles need to keep consistent. Thus, we propose a rotation-aware consistency (RAC) learning to handle this issue. As the activation patterns inevitable shift on the target domains and yield less discriminative RoIs (discussed in Sec. 1), we consider to pose RAC between the RoIs between the original samples and the style-hallucinated samples.
Given a typical two-stage oriented object detector, after extracting feature representation $\mathbf{F}^{4}$ from the backbone, the regional proposal network (RPN) generates $n$ horizontal region of interests (HRoIs) $\{\mathcal{H}_{i}\}$, where $i=1,\cdots,n$. Generally, $\mathcal{H}_{i}$ is under the format of $\{\overline{x},\overline{y},\overline{w},\overline{h}\}$, which describe the 2D location of a RoI. Then, $\{\mathcal{H}_{i}\}$ need to be mapped to rotated RoIs (RRoIs) $\{\mathcal{R}_{i}\}$ to represent the orientation information, where $\mathcal{R}_{i}$ is under the format of $(x,y,w,h,\theta)$.
An ideal domain generalized oriented object detector need to stabilize both rotation and content representation for both $\{\mathcal{H}_{i}\}$ and $\{\mathcal{R}_{i}\}$, despite the great style variation for samples from different domains.
The HRoIs’ feature representation of the original image $\mathbf{F}_{\mathcal{H}}$ and the style hallucinated image $\mathbf{\widetilde{F}}_{\mathcal{H}}$ are learnt from $\mathbf{F}$ and $\mathbf{\widetilde{F}}$ by region proposal and feature alignment, respectively. To maintain the content consistency before and after style hallucination, both HRoIs’ representations $\mathbf{F}_{\mathcal{H}}$ and $\mathbf{\widetilde{F}}_{\mathcal{H}}$ are projected into the embeddings $\mathbf{z}_{\mathcal{H}}\in\mathbb{R}^{d\times n}$ and $\mathbf{\widetilde{z}}_{\mathcal{H}}\in\mathbb{R}^{d\times n}$ by two MLPs $f_{1}(\cdot)$ and $f_{2}(\cdot)$, respectively. Here $d$ refers to the dimension size of each embedding.
Let $\mathbf{z}=[\mathbf{z}_{\mathcal{H}},\mathbf{\widetilde{z}}_{\mathcal{H}}]\in\mathbb{R}^{d\times 2n}$, and let $z_{j},z_{k}\in\mathbf{z}$, where $z_{j},z_{k}\in\mathbb{R}^{d\times 1}$, $j,k=1,\cdots,2n$ and $j\neq k$. Assume $z_{j,+}$ and $z_{j}$ construct a positive pair, the content consistency learning objective for HRoIs (denoted as $\mathcal{L}_{HCL}$) can be computed as
$\mathcal{L}_{HCL}=-\sum_{j=1}^{2n}\sigma(\mathcal{H}_{j})\cdot{\rm log}\frac{exp(z_{j}\cdot z_{j,+}/\tau)}{\sum_{j=1,j\neq k}^{2n}exp(z_{j}\cdot z_{k}/\tau)},$ (6)
where $\sigma(\cdot)$ is a binary function determined by the mean intersection over union (mIoU) value between a region of interest (e.g., $\mathcal{H}_{j}$) and ground truth (denoted as $\rm{gt}$), given by
$\sigma(\mathcal{H}_{j})=\begin{cases}1&\text{if mIoU$(\mathcal{H}_{j},\rm{gt})\geq$0.5 }\\ 0&\text{else}\end{cases}.$ (7)
[…]
To realize this objective, two MLPs $f_{3}(\cdot)$ and $f_{4}(\cdot)$ are used to project $\mathbf{F}_{\mathcal{R}}$ and $\mathbf{\widetilde{F}}_{\mathcal{R}}$ into embeddings, which we denote as $\mathbf{z}_{\mathcal{R}}\in\mathbb{R}^{d\times n}$ and $\mathbf{\widetilde{z}}_{\mathcal{R}}\in\mathbb{R}^{d\times n}$. Similarly, we have $\mathbf{z}^{\prime}=[\mathbf{z}_{\mathcal{R}},\mathbf{\widetilde{z}}_{\mathcal{R}}]\in\mathbb{R}^{d\times 2n}$, and let $z^{\prime}_{j},z^{\prime}_{k}\in\mathbf{z}^{\prime}$, where $z^{\prime}_{j},z^{\prime}_{k}\in\mathbb{R}^{d\times 1}$, $j,k=1,\cdots,2n$ and $j\neq k$. Assume $z^{\prime}_{j,+}$ and $z^{\prime}_{j}$ construct a positive pair, and the rotation-aware consistency learning objective for RRoIs (denoted as $\mathcal{L}_{RAC}$) can be computed as
$\mathcal{L}_{RAC}=-\sum_{j=1}^{2n}\sigma(\mathcal{R}_{j})\cdot{\rm log}\frac{exp(z^{\prime}_{j}\cdot z^{\prime}_{j,+}/\tau)}{\sum_{j=1,j\neq k}^{2n}exp(z^{\prime}_{j}\cdot z^{\prime}_{k}/\tau)}.$ (8)
3.4 Style Consistency Learning
For oriented objects, a domain generalized detector is supposed to show robust orientation prediction despite the great style variation. During its training, it is necessary to constraint the style hallucination to be non-geometry, so that the original samples and the style-hallucinated samples could carry the same structural information. We propose a style consistency (SEC) learning to realize this objective.
To begin with, two fully-connected layers $g_{1}(\cdot)$ and $g_{2}(\cdot)$ are used to map the oriented feature representation of original sample $\mathbf{F}_{\mathcal{R}}$ and style-hallucinated sample $\mathbf{\widetilde{F}}_{\mathcal{R}}$ to category-wise probability distribution $\mathbf{p}$ and $\mathbf{\widetilde{p}}$, given by
$\mathbf{p}=g_{1}(\mathbf{F}_{\mathcal{R}}),\quad\mathbf{\widetilde{p}}=g_{2}(\mathbf{\widetilde{F}}_{\mathcal{R}}).$ (9)
Both category-level scores $\mathbf{p}$ and $\mathbf{\widetilde{p}}$ need to be restricted in a way such that the pixel-level category remain unchanged in both original and style diversified samples. The Jensen-Shannon Divergence (JSD) is utilized to realize this style consistency objective, given by
$\mathcal{L}_{SEC}=\frac{1}{2}({\rm KL}[\mathbf{p}||\mathbf{\overline{p}}]+{\rm KL}[\mathbf{\widetilde{p}}||\mathbf{\overline{p}}]),$ (10)
where $\mathbf{\overline{p}}=(\mathbf{\widetilde{p}}+\mathbf{p})/2$, and ${\rm KL}$ denotes the Kullback-Leibler divergence.
To summarize, the proposed SEC poses a category-wise constraint between the original and style-hallucinated images, which allows the style-hallucinated images to perceive the same content of an aerial image.
3.5 Implementation Details
Assume a convention oriented object detector is optimized by the classification head $\mathcal{L}_{cls}$ and bounding box regression head $\mathcal{L}_{reg}$. Then, allowing for Eq. 68 and 10, the learning objective of GOOD is computed as
$\mathcal{L}=\mathcal{L}_{cls}+\mathcal{L}_{reg}+\mathcal{L}_{HC}+\mathcal{L}_{RAC}+\mathcal{L}_{SEC}.$ (11)
For classification $\mathcal{L}_{cls}$ and regression $\mathcal{L}_{reg}$, both original feature representation and style-hallucinated representation are involved for training.
Following the standard configuration of object detectors, ResNet-50 with Feature Pyramid Network (FPN) is used as the backbone. In FPN, each pyramid assigns 15 anchors per location. For the region proposal network, 1024 Region of Interests (RoIs) with a 1:3 positive to negative ratio are sampled for training. All the training schedules follow the default settings in mmdetection  [^1] https://github.com/open-mmlab/mmdetection.
The SGD optimizer is used for training with an initial learning rate of 0.01. The value of momentum and weight decay is set to be 0.9 and 0.0001, respectively. After each decay step, the learning rate is divided by 10. All the models are trained by a total of 20 epochs on a workstation. 4 V100 GPUs with a total batch size of 8 are used for training.
When inference, the CLIP-driven style hallucination is removed, and the detector directly predicts oriented bounding boxes on unseen target domains. 2,000 RoIs from each pyramid level are adapted before non-maximum suppression (NMS). A total of 2,000 RoIs are selected after NMS.
Setting Source Domain Unseen Target Domain #Cate. Categories
I FAIR1M DOTA-V1, DOTA-V1.5, DOTA-V2.0 10 airplane, ship, baseball field, tennis court, basketball court, football field, large vehicle, small vehicle, bridge, roundabout
II SODA DOTA-V1, DOTA-V1.5, DOTA-V2.0 7 airplane, ship, storage tank, large vehicle, small vehicle, helicopter, swimming pool
III DOTA-V1 SODA, FAIR1M 4 airplane, ship, large vehicle, small vehicle
IV DOTA-V1 SODA, FAIR1M, HRSC 1 ship
Table 1: Experimental settings for domain generalized oriented object detection. Following prior generalized driving-scene detection works [46, 29, 42], the shared object categories among different datasets are used to evaluate the domain generalized oriented object detector.
Method Source Domain: FAIR1M Source Domain: SODA Source Domain: DOTA
DOTA-1.0 DOTA-1.5 DOTA-2.0 DOTA-1.0 DOTA-1.5 DOTA-2.0 SODA FAIR1M
one-stage:
RetinaNet OBB [30] 28.78 28.31 26.34 37.82 37.72 35.08 39.39 29.65
FCOS OBB [41] 30.47 31.46 29.76 37.49 39.80 37.45 42.46 31.86
ATSS OBB [59] 35.89 31.12 29.39 37.38 38.30 36.41 41.12 34.67
R3Det [51] 35.10 34.79 33.79 38.77 39.33 37.39 44.11 36.51
SA2Net [15] 32.31 32.57 31.41 35.21 35.70 33.37 33.25 29.90
GWD [52] 38.59 39.56 37.35 38.70 39.77 38.62 46.46 40.31
KLD [53] 37.59 36.76 34.92 39.19 41.19 39.14 48.57 36.99
SASM [18] 35.69 35.83 33.82 41.20 42.28 40.15 47.34 38.15
RepPoint [26] 37.80 36.86 34.85 37.41 40.53 37.66 50.49 40.63
two-stage:
FR OBB [37] 33.15 33.17 30.79 34.98 36.89 34.89 44.14 34.46
MR OBB [17] 34.46 35.99 33.30 39.81 42.60 41.08 46.32 36.54
Oriented RCNN [48] 35.72 34.97 34.19 40.07 41.23 38.35 48.83 37.34
RoI-Tran [11] 36.82 37.02 35.21 41.93 43.19 40.92 52.06 41.18
ReDet [16] 40.46 41.06 38.35 42.18 44.25 41.05 54.85 43.09
GOOD (RoI-Tran) 37.94 37.55 36.31 44.03 45.86 42.14 54.11 42.74
$\uparrow$1.12 $\uparrow$0.53 $\uparrow$1.10 $\uparrow$2.10 $\uparrow$2.67 $\uparrow$1.22 $\uparrow$2.05 $\uparrow$1.56
GOOD (ReDet) 42.61 42.70 40.36 44.65 47.54 43.88 57.19 46.21
$\uparrow$2.15 $\uparrow$1.64 $\uparrow$2.01 $\uparrow$2.47 $\uparrow$3.29 $\uparrow$2.83 $\uparrow$3.08 $\uparrow$3.12
Table 2: Result comparison between the proposed GOOD and other oriented object detectors on unseen target domains. Evaluation metric $mAP_{50}$ (in %). For fair evaluation, all the compared methods use the ResNet-50 backbone pre-trained on ImageNet.
4 Experiment
4.1 Dataset & Evaluation Protocols
FAIR1M [39] is a fine-grained aerial object detection dataset with 37 corresponding fine-grained categories. It has more than 15,000 images and 1 million instances. To set up domain generalization setting which exists same object categories from cross-domain, we merge the 11 fine-grained airplane categories into a single airplane category, 9 fine-grained ship categories into a ship category, 10 fine-grained vehicle categories into both large vehicle and small vehicle categories. On the other hand, as the test set of FAIR1M is not publicly available, we randomly split the official training set as 8:2 as the source domain for training and target domain for inference, respectively.
DOTA [47] consists of 2,806 aerial images with 188,282 instances. The annotation of DOTA  [^2] https://captain-whu.github.io/DOTA/index.html has three versions, namely, DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0 [12]. DOTA-V1.0 has 15 object categories. DOTA-V1.5 and DOTA-V2.0 has one and three more object categories than DOTA-V1.0, respectively. Following the default setting of MMRotate  [^3] https://github.com/open-mmlab/mmrotate, the images are split into 1024$\times$1024 for training, and are merged for evaluation.
SODA [8] consists of 2,513 images with up to 872,069 instances from 9 classes. It is an oriented object detection benchmark for small objects. All the data split and pre-processing follow the default setting of MMRotate  [^4] https://github.com/shaunyuan22/SODA-mmrotate.
HRSC [31] is a ship detection dataset with arbitrary orientations. The training and testing split directly follows the default setting of MMRotate, where all the images are resized to 800$\times$800 for both training and testing.
Evaluation Protocols.
Following prior domain generalized driving-scene object detection protocols [46, 29, 42], the shared object categories among different datasets are used to evaluate the domain generalized object detector. The four domain generalization settings and the corresponding object categories are listed in Table 1.
Evaluation Metric. For all the experiments, Mean Average Precision (mAP) is used as the evaluation metric. In line with [47], mAP@0.5 is reported, which classifies a prediction as a true positive when it aligns with the ground-truth label and achieves an Intersection over Union (IOU) score of over 0.5 with the ground-truth bounding box.
4.2 Comparison with State-of-the-art
Method AP BF BG SV LV SP TC BC FF RA
R3Det [51] 59.4 34.1 10.2 16.6 4.0 65.5 68.7 22.4 42.5 27.8
RepPoint [26] 64.3 38.1 12.8 21.3 5.1 69.9 76.5 31.6 39.0 19.3
RoI-Tran [11] 60.6 41.0 13.6 12.3 5.6 76.0 62.3 29.3 45.3 22.2
ReDet [16] 77.9 34.6 12.5 10.6 7.6 76.8 77.1 41.7 40.4 25.2
GOOD (ReDet) 78.2 41.4 16.3 13.3 5.9 76.6 76.5 45.6 47.1 25.4
$\uparrow$0.3 $\uparrow$4.8 $\uparrow$3.8 $\uparrow$2.7 $\downarrow$1.7 $\downarrow$0.2 $\downarrow$0.6 $\uparrow$3.9 $\uparrow$6.7 $\uparrow$0.2
[…]
Method Target Domains
SODA FAIR1M HRSC
RoI-Tran [11] 67.20 35.05 35.94
GOOD (RoI-Tran) 69.92 39.26 38.87
$\uparrow$2.72 $\uparrow$4.21 $\uparrow$2.93
ReDet [16] 75.89 46.37 59.71
GOOD (ReDet) 78.97 48.44 62.08
$\uparrow$3.08 $\uparrow$2.07 $\uparrow$2.37
[…]
Component Target Domains
Style $\mathcal{L}_{HCL}$ $\mathcal{L}_{RCL}$ $\mathcal{L}_{SEC}$ D-1.0 D-1.5 D-2.0
42.18 44.25 41.05
$\checkmark$ 43.17 45.16 41.62
$\checkmark$ $\checkmark$ 43.24 45.77 42.08
$\checkmark$ $\checkmark$ $\checkmark$ 43.54 46.49 42.49
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 44.65 47.54 43.88
[…]
Method RMSD $\downarrow$ on Target Domains
SODA FAIR1M HRSC
RoI-Tran 0.137 0.106 0.084
GOOD (RoI-Tran) 0.091 0.058 0.043
$\downarrow$33.6$\downarrow$45.3$\downarrow$48.8%
ReDet 0.125 0.109 0.084
GOOD (ReDet) 0.086 0.049 0.038
$\downarrow$31.2$\downarrow$55.0$\downarrow$54.7%
[…]
The proposed GOOD along with extensive state-of-the-art oriented object detectors are involved for experimental evaluation. For fair comparison, all the detectors are trained using the ResNet-50 backbone with a pre-trained weight on ImageNet. Besides, under each domain generalization setting, each detector is implemented under the corresponding default parameter settings with the official source code.
(a) RoI-Trans
(b) GOOD (RoI-Trans)
(c) ReDet
(d) GOOD (ReDet)
Figure 5: T-sne visualization of the cross-domain feature space. Blue, green and orange denotes the target domains of SODA, FAIR1M, HRSC, respectively. The more uniformly distributed from these domains, the better feature generalization.
[…]
SODA as Source Domain The third column of Table 2 reports the results. The proposed GOOD yields an mAP of 44.65%, 47.54% and 43.88% on DOTA-1.0, -1.5 and -2.0, outperforming all the compared oriented object detectors.
DOTA-V1 as Source Domain The forth column of Table 2 reports the results. The proposed GOOD achieves the best generalization results, yielding an mAP of 57.19% and 46.21% on SODA and FAIR1M, respectively.
[…]
Generalized Ship Detection Table 4 reports the results when detecting one-type ship detection on unseen target domains. The proposed GOOD achieves the best results on all three unseen target domains, yielding an mAP of 78.97%, 48.44% and 62.08%, respectively.
4.3 Ablation Studies
On Each Component Table 5 studies the impact of each component in GOOD, namely, style hallucination, HRoI consistency learning ($\mathcal{L}_{HCL}$ in Eq. 6), rotation-aware consistency learning ($\mathcal{L}_{RCL}$ in Eq. 8) and style consistency learning ($\mathcal{L}_{SEC}$ in Eq. 10). Only implementing style hallucination slightly improves the generalization, yielding 0.99%, 0.91% and 0.57% mAP gain on DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0. Implementing consistency learning on both HRoIs ($\mathcal{L}_{HCL}$) and RRoIs ($\mathcal{L}_{RCL}$) can improve the generalization ability on unseen target domains. Especially, $\mathcal{L}_{RCL}$ yields a 0.30%, 0.72% and 0.41% mAP improvement on DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0 domain. Finally, the constraint on style hallucination ($\mathcal{L}_{SEC}$) can further improve the generalization.
Feature Space Visualization by t-SNE is implemented on both the proposed GOOD and its baseline model. Fig. 5 shows that the proposed GOOD allows the samples from source domain DOTA-V1 and the unseen target domain SODA, FAIR1M and HRSC to be more uniformly clustered than the baseline, which indicates its better generalization ability on unseen target domains.
[…]
On Style from Each Scale Table 7 systematically studies the impact of style hallucination on the feature representation from each scale, denoted as $\mathbf{F}^{1}$, $\mathbf{F}^{2}$, $\mathbf{F}^{3}$ and $\mathbf{F}^{4}$. The style hallucination on each scale positively contributes to the detector’s generalization ability.
Style Hallucination Target Domains
$\mathbf{F}^{1}$ $\mathbf{F}^{2}$ $\mathbf{F}^{3}$ $\mathbf{F}^{4}$ D-1.0 D-1.5 D-2.0
$\checkmark$ 43.25 46.12 42.53
$\checkmark$ $\checkmark$ 43.65 46.38 43.04
$\checkmark$ $\checkmark$ $\checkmark$ 43.96 46.83 43.56
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 44.65 47.54 43.88
[…]
On Different Distance Metric Table 8 investigates the impact of different distance metric on $\mathcal{L}_{SC}$ (in Eq. 10). The JSD distance metric in $\mathcal{L}_{SEC}$ is compared with four commonly-used distance metrics, namely, $l$-2 distance and K-L distance. The style consistency learning from $l$-2 and K-L distance is less significant than the JSD distance.
Distance Metric Target Domains
SODA FAIR1M
$l$-2 loss 56.13 45.56
K-L loss 56.52 46.07
JSD loss 57.19 46.21
[…]
4.4 Visualization
Fig. 6 visualizes some detection results on unseen target domains from the ReDet baseline and the proposed GOOD. The proposed GOOD shows more reasonable and more precise rotated bounding boxes on these unseen target domains. More visualization results in the supplementary materials.
5 Conclusion
In this paper, we make an early exploration to learn domain generalized oriented object detection from a single source domain. It is particularly challenging as the detector need to represent not only stable content but also precise orientation under the cross-domain style variation. In the proposed GOOD, the CLIP-driven style hallucination is constrainted by the proposed rotation-aware content consistency learning and style-consistency learning, which allows the model to learn generalizable rotation and content representation.
Extensive experiments on a series of domain generalization settings show that the proposed GOOD has significantly stronger generalization than existing state-of-the-art oriented object detectors.
Limitation & Future Work. We focus on existing domain generalized detection settings, where the cross-domain object categories are consistent. In the future, we will GOOD under open-world scenarios, which can further discover unknown categories from different domains.